# FLARE: Agile Flights for Quadrotor Cable-Suspended Payload System via Reinforcement Learning

### v1

@@ III-B Payload targeting
As shown in Fig. LABEL:genesis_task2 and Fig. LABEL:task2_method2, our method successfully navigates to a sequence of 10 payload targets, forming a double-loop, star-shaped trajectory. Additionally, we compare the velocity performance and completion time with those of scenario 1, where the identical targets are treated as quadrotor waypoints, as summarized in Table III.
TABLE III: Performance Metrics Comparison on 10 Identical Targets (0.2m Threshold)
Scenario Max Vel. Avg Vel. Time
(m/s) (m/s) (s)
1 $4.05$ $2.55$ $8.57$
2 $3.33$ $1.78$ $8.15$
The aforementioned comparison of the policies from scenario 1 and scenario 2 reveals that the latter learns to exploit the natural swing characteristics of the cable. This learned behavior allows the payload to reach targets by harnessing its momentum, resulting in a more efficient flight with a shorter travel distance and duration.
@@ III-C Gate traversal
TABLE IV: Performance Metrics Comparison Over 8 Simulation Tracks: Maximum Velocity, Traversal Time, Computational Time, Quadrotor Traversal Deviation, Payload Traversal Deviation, and Success Rate.
@@
We validate our method’s ability by performing both one-time and sequential gate traversal, as shown in Fig. 6. For performance evaluation, we benchmark our approach against Impactor [12] in the one-time traversal setting. This comparison is conducted across 8 trials, where the final waypoint is positioned within the range of [-2.0, 2.0] m. We evaluate both methods based on 6 important metrics as presented in TABLE LABEL:sim_task3_tabular. To offer a more detailed view, Fig. 7 illustrates the quadrotor and payload trajectories of a representative track governed by these two methods.
Figure 7: Time-series comparison of the trajectories for our method versus Impactor [12] on a representative gate traversal track. The shaded regions highlight the traversal phase (blue: our method; orange: Impactor), defined as the duration from the system’s initial entry to its full exit of the gate.
@@ IV-1 Agile waypoint passing
We design three distinct tracks to evaluate the performance of the learned policy in scenario 1, as one representative trajectory shown in Fig. 1(a). The 3D trajectories of both the quadrotor and the payload, presented in Fig. 9, demonstrate that the policy can effectively tackle diverse tracks in the waypoint-passing scenario.
@@
TABLE V: Performance Comparison of Simulation and Real-World Experiments on 3 Different Tracks
@@
Track 1
(4 waypoints)
Simulation 1.33 4.71
@@
Track 2
(7 waypoints)
Simulation 2.36 5.40
@@
Track 3
(14 waypoints)
Simulation 6.00 4.72
@@ IV-2 Payload targeting
The performance of the learned policy in scenario 2 is assessed by its ability to reach a given target at a distance of 1.5 m, as shown in Fig. 1(b). The motion data recorded in several trials shows that the policy achieves an average payload targeting error of 0.12 m. The payload’s average and maximum velocities are 1.61 m/s and 3.74 m/s, respectively. The quadrotor’s average and maximum velocities are 1.42 m/s and 2.98 m/s, respectively. These results indicate that the system exhibits high targeting accuracy and excellent speed performance in this scenario.
@@ IV-3 Gate traversal
The system’s gate traversal capability in scenario 3 is assessed by requiring it to navigate through both a single gate and two consecutive gates, as shown in Fig. 1(c). Notably, this scenario is particularly demanding as the combined size of the quadrotor and swinging payload exceeds the gate’s diameter, necessitating agile maneuvers, as the specific configuration detailed in Fig. 8.
In this scenario, both the quadrotor and the payload successfully navigate through narrow gates. During the consecutive traversal, the deviations from the center of the first gate are 0.21 m (quadrotor) and 0.17 m (payload), followed by deviations of 0.25 m and 0.26 m at the second. The entire maneuver is completed in a swift 1.57 s, with the quadrotor achieving an average velocity of 2.19 m/s and a maximum velocity of 5.72 m/s. These results underscore our policy’s remarkable aenility and safety.

### v2

@@ III-B Payload targeting
As shown in Fig. LABEL:task2_method2, our method successfully completes a sequence of 10 payload targets, forming a double-loop, star-shaped trajectory. Additionally, we compare the velocity performance and completion time with those of scenario 1, where the identical targets are treated as quadrotor waypoints, as summarized in Fig. LABEL:genesis_task2 and Table III.
(a) Visualized quadrotor and payload trajectories (2 laps, 0.2m threshold) on the star-shaped track governed by different policies.
The trajectory comparison presented in Fig. 6(a) illustrates that the policy from scenario 2 yields a more compact trajectory. This compactness reveals that the policy guides the quadrotor to brake earlier, leveraging the cable’s natural swing dynamics to reach the targets. This is further confirmed by Table III, which demonstrates that the policy from scenario 2 completes the track with reduced total quadrotor flight distance, validating corresponding superior efficiency.
TABLE III: Performance Metrics Comparison on 10 Identical Targets in Scenario 2 (0.2m Threshold)
Scenario Max Vel. Avg Vel. Time Quad. Dist.
(m/s) (m/s) (s) (m)
1 $4.05$ $2.55$ $8.57$ 21.85
2 $3.33$ $1.78$ $8.15$ 15.72
@@ III-C Gate traversal
TABLE IV: Performance Metrics Comparison Over 8 Simulation Trials in Scenario 3: Maximum Velocity, Traversal Time, Computational Time, Quadrotor Traversal Deviation, Payload Traversal Deviation, and Success Rate.
@@
Figure 7: Time-series comparison of the trajectories for our method versus Impactor [12] in scenario 3 (gate traversal). The shaded regions highlight the traversal phase (blue: our method; orange: Impactor), defined as the duration from the system’s initial entry to its full exit of the gate.
We validate our method’s ability by performing both one-time and sequential gate traversal, as shown in Fig. 6. For performance evaluation, we benchmark our approach against the aforementioned state-of-the-art optimization-based Impactor [12] in the one-time traversal setting. While effective for mode-switching planning, Impactor suffers from high computational costs, preventing its online deployment. This comparison is conducted across 8 trials, where the final waypoint is positioned within the range of [-2.0, 2.0] m. We evaluate both methods based on 6 important metrics as presented in TABLE LABEL:sim_task3_tabular. To offer a more detailed view, Fig. 7 illustrates the quadrotor and payload trajectories of a representative track governed by these two methods.
@@ IV-1 Agile waypoint passing
We design three distinct tracks to evaluate the performance of the learned policy in scenario 1, as one representative trajectory shown in Fig. LABEL:real-exp-fig(a). The 3D trajectories of both the quadrotor and the payload, presented in Fig. 9, demonstrate that the policy can effectively tackle diverse tracks in the waypoint-passing scenario.
@@
TABLE V: Performance Comparison of Simulation and Real-World Experiments on 3 Different Tracks in Scenario 1.
@@
Track 1 (4 waypoints) Simulation 1.33 4.71
@@
Track 2 (7 waypoints) Simulation 2.36 5.40
@@
Track 3 (14 waypoints) Simulation 6.00 4.72
@@ IV-2 Payload targeting
The performance of the learned policy in scenario 2 is assessed by its ability to reach a given target at a distance of 1.5 m, as shown in Fig. LABEL:real-exp-fig(b). The motion data recorded in several trials shows that the policy achieves an average payload targeting error of 0.12 m. The payload’s average and maximum velocities are 1.61 m/s and 3.74 m/s, respectively. The quadrotor’s average and maximum velocities are 1.42 m/s and 2.98 m/s, respectively. These results indicate that the system exhibits high targeting accuracy and excellent speed performance in this scenario.
@@ IV-3 Gate traversal
The system’s gate traversal capability in scenario 3 is assessed by requiring it to navigate through both a single gate and two consecutive gates, as shown in Fig. LABEL:real-exp-fig(c). Notably, this scenario is particularly demanding as the combined size of the quadrotor and swinging payload exceeds the gate’s diameter, necessitating agile maneuvers, as the specific configuration detailed in Fig. 8.
In this scenario, both the quadrotor and the payload successfully navigate through narrow gates. During the consecutive traversal, the deviations from the center of the first gate are 0.21 m (quadrotor) and 0.17 m (payload), followed by deviations of 0.25 m and 0.26 m at the second. The entire maneuver is completed in a swift 1.57 s, with the quadrotor achieving an average velocity of 2.19 m/s and a maximum velocity of 5.72 m/s. These results underscore our policy’s remarkable agility and safety.